# LiteASR: Efficient Automatic Speech Recognition with Low-Rank Approximation

### v1

@@ Abstract
Modern automatic speech recognition (ASR) models, such as OpenAI’s Whisper, rely on deep encoder-decoder architectures, and their encoders are a critical bottleneck for efficient deployment due to high computational intensity. We introduce LiteASR, a low-rank compression scheme for ASR encoders that significantly reduces inference costs while maintaining transcription accuracy. Our approach leverages the strong low-rank properties observed in intermediate activations: by applying principal component analysis (PCA) with a small calibration dataset, we approximate linear transformations with a chain of low-rank matrix multiplications, and further optimize self-attention to work in the reduced dimension. Evaluation results show that our method can compress Whisper large-v3’s encoder size by over 50%, matching Whisper medium’s size with better transcription accuracy, thereby establishing a new Pareto-optimal frontier of efficiency and performance. The code of LiteASR is available at https://github.com/efeslab/LiteASR.
@@
Keisuke Kamahori1,2  Jungo Kasai2  Noriyuki Kojima2  Baris Kasikci1 1University of Washington 2Kotoba Technologies Inc. {kamahori,baris}@cs.washington.edu, {jkasai,nkojima}@kotoba.tech
@@ 1 Introduction
Figure 1: The relationship between encoder size and accuracy, as measured by word error rate (WER), for models in the Whisper family. The stars denote variants compressed via our method, which achieves the Pareto-optimal balance between accuracy and efficiency.
@@
In this work, we propose LiteASR, a novel compression scheme that targets ASR encoders by exploiting the low-rank structure of hidden activations during inference. A key insight driving our approach is that intermediate activations, both in self-attention and multi-layer perception (MLP) layers, consistently exhibit low-rank properties across a wide variety of inputs. This phenomenon stems from ASR encoders’ use of Mel spectrograms, the 2D time-frequency audio representations. Real-world audio (e.g., human speech) exhibits strong correlations between frequency components Huang et al. (2012); Zergat and Amrouche (2013); Tian et al. (2024); Kacha et al. (2020), resulting in low-rank characteristics of the intermediate features.
Our method first analyzes the low-rank properties of activations using a small amount of calibration data. We then perform a principal component analysis (PCA) Wold et al. (1987) to extract the dominant components and approximate linear transformations with rank-$k$ projections. This factorization allows each weight matrix to be expressed as the product of two lower-rank matrices, thereby reducing the total number of floating point operations (FLOPs) required for inference. We employ an adaptive mechanism based on the threshold to determine the optimal degree of low-rank approximation for each layer.
To further capitalize on the optimization, we also modify the self-attention algorithm to operate in the reduced dimension. We implement a specialized GPU kernel based on FlashAttention Dao et al. (2022) to accelerate the computation of attention scores and outputs.
Our evaluation shows that LiteASR achieves a Pareto-optimal balance between speed and accuracy (see Figure 1). When applied to Whisper large-v3, LiteASR reduces the encoder size by approximately 40%, yielding an execution speedup of around 1.4x with negligible accuracy loss. In alternative configurations, we further reduce the model size to less than half, resulting in a model comparable in size to Whisper medium, while delivering improved accuracy. We also demonstrate the applicability of the method across different languages and models (§4).
@@
    We introduce LiteASR, a compression method for ASR encoders using a low-rank approximation of activation values. This method approximates linear layers with a chain of low-rank matrix multiplications and optimizes self-attention to operate in a reduced dimension.
@@
    We present a comprehensive evaluation demonstrating that our method achieves a Pareto-optimal balance between accuracy and efficiency.
@@ 2.2 Compute Requirements of ASR Models
Since the encoder often processes long sequences (e.g., fixed at 1500 for Whisper), it often emerges as the primary runtime bottleneck. Figure 2 shows the latency breakdown between the encoder and decoder across three hardware setups (NVIDIA RTX 4090, NVIDIA RTX A6000, and Apple M1 Pro), two models (Whisper large-v3 and Whisper large-v3-turbo), and two batch sizes (1 and 8) [^1] We use vLLM Kwon et al. (2023) (ver. 0.7.0) and MLX Hannun et al. (2023) (ver. 0.21.1) to transcribe a sample audio clip from the ESB dataset Gandhi et al. (2022)..
@@ 3 Methodology
Our method, LiteASR, compresses the ASR encoder by extracting the low-rank features from activations at different layers of the model. To do so, we first use calibration data to analyze activations and then convert the dense matrix multiplication within the model to the product of low-rank matrices (Figure 3 shows a simplified overview of the method). We further modify the self-attention algorithm to work efficiently on reduced dimensions. In this section, we explain the methodologies in detail.
@@ 3.1 Analyzing Activations in Transformers
where the weight matrix $W\in\mathbb{R}^{D_{\text{in}}\times D_{\text{out}}}$ and the bias vector $b\in\mathbb{R}^{D_{\text{out}}}$ are learnable model parameters. Here, the input activations $X\in\mathbb{R}^{L\times D_{\text{in}}}$ produce the output activations $Y\in\mathbb{R}^{L\times D_{\text{out}}}$ during the forward pass. In this notation, $D_{\text{in}}$ and $D_{\text{out}}$ denote the input and output dimensions of the layer, respectively, and $L$ is the sequence length [^2] For Whisper encoders, this is always $1500$..
@@ 3.2.1 How to Choose $k$
To maximize the GPU efficiency, we further restrict $k$ to be a multiple of $16$. Therefore, we choose $k$ as the smallest multiple of $16$ that satisfies both of Equation 6 and Equation 7. We empirically find that $\theta$ values between $0.99$ and $0.999$ achieve a good balance between accuracy and efficiency. A detailed sensitivity study on the choice of $\theta$ is provided in §4.
@@ 3.2.2 Optimizing Self-Attention
The standard approach (Equation 10) computes $Q_{i}$ and $K_{i}^{\top}$ separately and then multiplies them, which requires approximately $\mathcal{O}(L^{2}D_{\text{head}})$ FLOPs. In contrast, Equation 12 allows us to first compute the smaller matrix product $W_{Q_{2}}^{i}W_{K_{2}}^{i\,\top}$ and then multiply by $A$ and $B$, reducing the computational cost to $\mathcal{O}(L\,k_{Q}\,k_{K}+L^{2}\,\min(k_{Q},k_{K}))$. This is beneficial when $\min(k_{Q},k_{K})<D_{\text{head}}$ [^3] We take the minimum of $k_{Q}$ and $k_{K}$ because we can choose the multiplication order to minimize computation.. Thus, we adopt Equation 12 if the rank is sufficiently small.
@@
Moreover, since each row of $S_{i}$ sums to $1$, the bias term simplifies [^4] Here, $b_{V}^{i}$ is broadcasted across each row of $S_{i}$.:
@@ 4.1 Setup
Regarding efficiency, we evaluated the encoder latency on NVIDIA RTX 4090, NVIDIA RTX A6000, and Apple M1 Pro. For GPUs, we modify OpenAI’s Whisper implementation [^5] https://github.com/openai/whisper to use CUDA Graph with PyTorch Ansel et al. (2024) (ver. 2.5.1), and we use Triton Tillet et al. (2019) (ver. 3.2.0) for a customized self-attention GPU kernel. On the Apple device, we use MLX Hannun et al. (2023) (ver. 0.21.1). The presented latency data are averaged over 10 runs. Note that the encoder always takes fixed-length audio as input, so the computation requirement is exactly the same for different data.
@@ 4.2 Accuracy Evaluation
Table 1 compares the WER and encoder size. LiteASR is evaluated on Whisper large-v3 and Whisper large-v3-turbo models, with Whisper medium as a reference. The quality-focused configuration (a) cuts model size by over 30% with less than 0.1 percentage points degradation of WER for both Whisper large-v3 and Whisper large-v3-turbo. For more efficiency-focused scenarios, configuration (b) reduces encoder size by over 40% with comparable WER for Whisper large-v3, and about 2.5 points degradation for Whisper large-v3-turbo. Configuration (c) compresses Whisper large-v3 model to less than half, matching Whisper medium’s size, with better WER by about 3.5 points. Overall, LiteASR significantly reduces the model size while largely maintaining accuracy.
@@ 4.3 Efficiency Evaluation
Figure 4 presents the efficiency evaluation results, measuring the speedup of end-to-end latency of the encoder execution compared to the original model. LiteASR consistently achieves speed improvements across all three hardware setups, with average speedups of 1.29x for (a), 1.38x for (b), and 1.54x for (c). The best performance is observed with the RTX 4090 using (c), reaching a peak speedup of 1.57x.
@@ 4.4.1 Compression Ratio per Layer
Figure 6 illustrates the compression ratio (i.e., defined as the quotient of $k$ divided by the original dimension size) for each linear layer within the Whisper large-v3 encoder. The data is presented for configuration (c). In general, the initial layers allow for more substantial compression, with some exceptions, such as the FC2 stage in layer 21. This tendency is most pronounced in FC2 layers, where the earlier layers exhibit a compression ratio of less than 0.2, whereas the subsequent layers reach values larger than 0.8. Among the layers, the Q/K projection and FC1 layers display a smaller compression ratio compared to other layers.
@@ 4.4.3 Sensitivity to Languages
To further investigate how LiteASR generalizes to out-of-distribution data and its sensitivity to languages, we extend our evaluation to non-English benchmarks. We use MLS Pratap et al. (2020) for French and German, and the JSUT basic5000 Sonobe et al. (2017) for Japanese [^6] For Japanese, we use the character error ratio (CER) instead of the WER since Japanese does not have explicit word boundaries.. Here, we use the same English calibration data as in previous experiments to compress Whisper large-v3, and evaluate its accuracy on non-English audio. The results presented in Table 2 demonstrate LiteASR’s robustness: for (a), there is almost no degradation in accuracy, and even for (c), the degradation is less than 2 percentage points in WER/CER. In some cases, such as with German, we even observe an improvement in accuracy.
@@ 4.4.4 Sensitivity to Models
We also evaluate on Canary 1B Puvvada et al. (2024), NVIDIA’s state-of-the-art ASR model, to determine LiteASR’s applicability to a broader range of models. The encoder of Canary employs the FastConformer architecture Rekesh et al. (2023), and our optimizations are confined to linear layers within the feed-forward and self-attention modules, leaving the convolution modules unaltered. Table 3 presents the encoder size and the average WER for ESB datasets. The data indicates that there is minimal degradation in the WER, although the reduction in size is moderate compared to the Whisper models, achieving approximately a 10% reduction for configuration (c).
@@ 5.1 Efficient ASR Inference
Several prior works have aimed to enhance ASR model efficiency. FasterWhisper uses optimized inference kernels SYSTRAN (2023), while WhisperX further improves it for long-form audio Bain et al. (2023). Whisper.cpp is a C/C++ implementation for portability on both the CPU and GPU Gerganov (2023). Whisper_streaming supports live transcription for streaming purposes Macháček et al. (2023). NVIDIA’s NeMo is a modular toolkit for deploying speech models Harper et al. . However, they do not effectively reduce ASR encoder computational demands. Some works provide model weight quantization, but they are limited to weights (weight-only quantization) and do not accelerate the compute-bound encoder inference. Our approach can be integrated with these frameworks.
Various studies, including Whisper large-v3-turbo, Distill-Whisper, and Kotoba-Whisper use distillation techniques to shrink decoder size Radford et al. (2023); Gandhi et al. (2023); Kotoba Technologies (2024). Other approaches combine distillation with quantization or lightweight modular ASR fine-tuning for underrepresented languages Shao et al. (2023); Ferraz et al. (2023). Our work complements these efforts by further reducing the encoder’s computational requirements.
@@ 5.2 Model Compression with Low-Rank Approximation
The low-rank approximation has been used to compress machine learning models, such as for parameter-efficient fine-tuning Hu et al. (2021) or the LLM’s KV cache compression Liu et al. (2024); Chang et al. (2024). Yu and Wu (2023) has suggested that activations in Transformer models exhibit low-rank and compressed models, mainly targeting vision models. However, their method is limited to linear layers, leaving self-attention layers unoptimized, and its applicability to speech models has not been studied.
@@ 6 Conclusion
In this work, we introduced a compression method for ASR encoders that leverages the inherent low-rank structure of activations in linear layers. By applying the PCA algorithm, this method approximates linear layers with a chain of low-rank matrix multiplications and optimizes self-attention to operate in a reduced dimension. Our comprehensive evaluation demonstrates that our method achieves a Pareto-optimal balance between accuracy and efficiency, paving the way for more efficient ASR deployments for both on-device and data center environments.

### v2

@@ Abstract
Modern automatic speech recognition (ASR) models, such as OpenAI’s Whisper, rely on deep encoder-decoder architectures, and their encoders are a critical bottleneck for efficient deployment due to high computational intensity. We introduce LiteASR, a low-rank compression scheme for ASR encoders that significantly reduces inference costs while maintaining transcription accuracy. Our approach leverages the strong low-rank properties observed in intermediate activations: by applying principal component analysis (PCA) with a small calibration dataset, we approximate linear transformations with a chain of low-rank matrix multiplications, and further optimize self-attention to work in reduced dimensionality. Evaluation results show that our method can compress Whisper large-v3’s encoder size by over 50%, matching Whisper medium’s size with better transcription accuracy, thereby establishing a new Pareto frontier of accuracy and efficiency. The code of LiteASR is available at https://github.com/efeslab/LiteASR.
@@
Keisuke Kamahori1,2  Jungo Kasai2  Noriyuki Kojima2  Baris Kasikci1 1University of Washington 2Kotoba Technologies, Inc. {kamahori,baris}@cs.washington.edu, {jkasai,nkojima}@kotoba.tech
@@ 1 Introduction
Figure 1: The relationship between encoder size and accuracy, as measured by word error rate (WER), for models in the Whisper family. The stars denote variants compressed via our method, which achieves an optimal trade-off between accuracy and efficiency.
@@
In this work, we propose LiteASR, a novel compression scheme that targets ASR encoders by exploiting the low-rank structure of hidden activations during inference. A key insight driving our approach is that intermediate activations, both in self-attention and multi-layer perceptron (MLP) layers, consistently exhibit low-rank properties across a wide variety of inputs. This phenomenon stems from ASR encoders’ use of Mel spectrograms, the 2D time-frequency audio representations. Real-world audio (e.g., human speech) exhibits strong correlations between frequency components Huang et al. (2012); Zergat and Amrouche (2013); Tian et al. (2024); Kacha et al. (2020), resulting in low-rank characteristics of the intermediate features.
Our method first analyzes the low-rank properties of activations using a small amount of calibration data. We then perform a principal component analysis (PCA) Wold et al. (1987) to extract the dominant components and approximate linear transformations with rank-$k$ projections. This factorization allows each weight matrix to be expressed as the product of two lower-rank matrices, thereby reducing the total number of floating-point operations (FLOPs) required for inference. We employ an adaptive mechanism based on the threshold to determine the optimal degree of low-rank approximation for each layer.
To further capitalize on the optimization, we also modify the self-attention algorithm to operate in reduced dimensionality. We implement a specialized GPU kernel based on FlashAttention Dao et al. (2022) to accelerate the computation of attention scores and outputs.
Our evaluation shows that LiteASR achieves an optimal trade-off between accuracy and efficiency (see Figure 1). When applied to Whisper large-v3, LiteASR reduces the encoder size by approximately 40%, yielding an execution speedup of around 1.4x with negligible accuracy loss. In alternative configurations, we further reduce the model size to less than half, resulting in a model comparable in size to Whisper medium, while delivering improved accuracy. We also demonstrate the applicability of the method across different languages and models (§4).
@@
    We introduce LiteASR, a compression method for ASR encoders using a low-rank approximation of activation values. This method approximates linear layers with a chain of low-rank matrix multiplications and optimizes self-attention to operate in reduced dimensionality.
@@
    We present a comprehensive evaluation demonstrating that our method achieves a Pareto frontier of accuracy and efficiency.
@@ 2.2 Compute Requirements of ASR Models
Since the encoder often processes long sequences (e.g., fixed at 1500 for Whisper), it often emerges as the primary runtime bottleneck. Figure 2 shows the latency breakdown between the encoder and decoder across three hardware setups (NVIDIA RTX 4090, NVIDIA RTX A6000, and Apple M1 Pro), two models (Whisper large-v3 and Whisper large-v3-turbo), and two batch sizes (1 and 8). [^1] We use vLLM Kwon et al. (2023) (ver. 0.7.0) and MLX Hannun et al. (2023) (ver. 0.21.1) to transcribe a sample audio clip from the ESB dataset Gandhi et al. (2022).
@@ 3 Methodology
Our method, LiteASR, compresses the ASR encoder by extracting the low-rank features from activations at different layers of the model. To do so, we first use calibration data to analyze activations and then convert the dense matrix multiplication within the model to the product of low-rank matrices (Figure 3 shows a simplified overview of the method). We further modify the self-attention algorithm to work efficiently in reduced dimensionality. In this section, we explain the methodologies in detail.
@@ 3.1 Analyzing Activations in Transformers
where the weight matrix $W\in\mathbb{R}^{D_{\text{in}}\times D_{\text{out}}}$ and the bias vector $b\in\mathbb{R}^{D_{\text{out}}}$ are learnable model parameters. Here, the input activations $X\in\mathbb{R}^{L\times D_{\text{in}}}$ produce the output activations $Y\in\mathbb{R}^{L\times D_{\text{out}}}$ during the forward pass. In this notation, $D_{\text{in}}$ and $D_{\text{out}}$ denote the input and output dimensions of the layer, respectively, and $L$ is the sequence length. [^2] For Whisper encoders, this is always $1500$.
@@ 3.2.1 How to Choose $k$
To maximize the GPU efficiency, we further restrict $k$ to be a multiple of $16$. Therefore, we choose $k$ as the smallest multiple of $16$ that satisfies both Equation 6 and 7. We empirically find that $\theta$ values between $0.99$ and $0.999$ achieve a good balance between accuracy and efficiency. A detailed sensitivity study on the choice of $\theta$ is provided in §4.
@@ 3.2.2 Optimizing Self-Attention
The standard approach (Equation 10) computes $Q_{i}$ and $K_{i}^{\top}$ separately and then multiplies them, which requires approximately $\mathcal{O}(L^{2}D_{\text{head}})$ FLOPs. In contrast, Equation 12 allows us to first compute the smaller matrix product $W_{Q_{2}}^{i}W_{K_{2}}^{i\,\top}$ and then multiply by $A$ and $B$, reducing the computational cost to $\mathcal{O}(L\,k_{Q}\,k_{K}+L^{2}\,\min(k_{Q},k_{K}))$. This is beneficial when $\min(k_{Q},k_{K})<D_{\text{head}}$. [^3] We take the minimum of $k_{Q}$ and $k_{K}$ because we can choose the multiplication order to minimize computation. Thus, we adopt Equation 12 if the rank is sufficiently small.
@@
Moreover, since each row of $S_{i}$ sums to $1$, the bias term simplifies: [^4] Here, $b_{V}^{i}$ is broadcast across each row of $S_{i}$.
@@ 4.1 Setup
Regarding efficiency, we evaluate the encoder latency on NVIDIA RTX 4090, NVIDIA RTX A6000, and Apple M1 Pro. For GPUs, we modify OpenAI’s Whisper implementation [^5] https://github.com/openai/whisper to use CUDA Graph with PyTorch Ansel et al. (2024) (ver. 2.5.1), and we use Triton Tillet et al. (2019) (ver. 3.2.0) for a customized self-attention GPU kernel. On the Apple device, we use MLX Hannun et al. (2023) (ver. 0.21.1). The presented latency data are averaged over 10 runs. Note that the encoder always takes fixed-length audio as input, so the computational requirements are exactly the same for different data.
@@ 4.2 Accuracy Evaluation
Table 1 compares the WER and encoder size. LiteASR is evaluated on Whisper large-v3 and Whisper large-v3-turbo models, with Whisper medium as a reference. The quality-focused configuration (a) cuts model size by over 30% with an increase of less than 0.1 percentage points in WER for both Whisper large-v3 and Whisper large-v3-turbo. For more efficiency-focused scenarios, configuration (b) reduces encoder size by over 40% with comparable WER for Whisper large-v3, and about 2.5 points degradation for Whisper large-v3-turbo. Configuration (c) compresses Whisper large-v3 model to less than half, matching Whisper medium’s size, with better WER by about 3.5 points. Overall, LiteASR significantly reduces the model size while largely maintaining accuracy. We emphasize that, unlike typical knowledge distillation methods which require substantial data and compute, [^6] For example, Distill-Whisper Gandhi et al. (2023) used ¿20k audio hours and the distillation took days on TPU v4-8. our approach achieves significant compression without accuracy degradation using only 100 audio clips and under 10 minutes on a single GPU.
@@ 4.3 Efficiency Evaluation
Figure 4 presents the efficiency evaluation results, measuring the speedup of end-to-end latency of the encoder execution compared to the original model. LiteASR consistently achieves latency improvements across all three hardware setups, with average speedups of 1.29x for (a), 1.38x for (b), and 1.54x for (c). The best performance is observed with the RTX 4090 using (c), reaching a peak speedup of 1.57x.
@@ 4.4.1 Compression Ratio per Layer
Figure 6 illustrates the compression ratio (i.e., defined as the quotient of $k$ divided by the original dimension size) for each linear layer within the Whisper large-v3 encoder. The data are presented for configuration (c). In general, the initial layers allow for more substantial compression, with some exceptions, such as the FC2 stage in layer 21. This tendency is most pronounced in FC2 layers, where the earlier layers exhibit a compression ratio of less than 0.2, whereas the subsequent layers reach values larger than 0.8. Among the layers, the Q/K projection and FC1 layers display a smaller compression ratio compared to other layers.
@@ 4.4.3 Sensitivity to Languages
To further investigate how LiteASR generalizes to out-of-distribution data and its sensitivity to languages, we extend our evaluation to non-English benchmarks. We use MLS Pratap et al. (2020) for French and German, and the JSUT basic5000 Sonobe et al. (2017) for Japanese. [^7] For Japanese, we use the character error rate (CER) instead of the WER since Japanese does not have explicit word boundaries. Here, we use the same English calibration data as in previous experiments to compress Whisper large-v3, and evaluate its accuracy on non-English audio. The results presented in Table 2 demonstrate LiteASR’s robustness: for (a), there is almost no degradation in accuracy, and even for (c), the degradation is less than 2 percentage points in WER/CER. In some cases, such as with German, we even observe an improvement in accuracy.
@@ 4.4.4 Sensitivity to Models
We also evaluate on Canary 1B Puvvada et al. (2024), NVIDIA’s state-of-the-art ASR model, to determine LiteASR’s applicability to a broader range of models. The encoder of Canary employs the FastConformer architecture Rekesh et al. (2023), and our optimizations are confined to linear layers within the feed-forward and self-attention modules, leaving the convolution modules unaltered. Table 3 presents the encoder size and the average WER for ESB datasets. The data indicate that there is minimal degradation in the WER, although the reduction in size is moderate compared to the Whisper models, achieving approximately a 10% reduction for configuration (c).
@@ 5.1 Efficient ASR Inference
Several prior works have aimed to enhance the efficiency of ASR models. FasterWhisper uses optimized inference kernels SYSTRAN (2023), while WhisperX further improves it for long-form audio Bain et al. (2023). Whisper.cpp is a C/C++ implementation for portability on both the CPU and GPU Gerganov (2023). Whisper_streaming supports live transcription for streaming purposes Macháček et al. (2023). NVIDIA’s NeMo is a modular toolkit for deploying speech models Harper et al. . However, they do not effectively reduce ASR encoder computational demands. Some works provide model weight quantization, but they are limited to weights (weight-only quantization) and do not accelerate the compute-bound encoder inference. Our approach can be integrated with these frameworks.
Various studies, including Whisper large-v3-turbo, Distill-Whisper, and Kotoba-Whisper use distillation techniques to shrink decoder size Radford et al. (2023); Gandhi et al. (2023); Kotoba Technologies (2024). Other approaches combine distillation with quantization or lightweight modular ASR fine-tuning for underrepresented languages Shao et al. (2023); Ferraz et al. (2023). Our work complements these efforts by further reducing the encoder’s computational requirements. Unlike these methods that require substantial data and compute resources, often tailored for specific downstream tasks, our work incurs minimal data/compute overhead. Our work also complements them by further reducing the encoder’s computational requirements; for instance, our compression technique can be effectively applied to models like Whisper large-v3-turbo, which is derived from the larger Whisper large-v3 model.
@@ 5.2 Model Compression with Low-Rank Approximation
The low-rank approximation has been used to compress machine learning models, such as for parameter-efficient fine-tuning Hu et al. (2021) or the LLM’s KV cache compression Liu et al. (2024); Chang et al. (2024). Yu and Wu (2023) have suggested that activations in Transformer models exhibit low-rank and compressed models, mainly targeting vision models. However, their method is limited to linear layers, leaving self-attention layers unoptimized, and its applicability to speech models has not been studied. Another work by Winata et al. (2020) presented a model architecture that adopts low-rank weights at train-time, whereas we propose a post-training technique for compressing pre-trained ASR models by analyzing activation patterns, avoiding costly retraining.
@@ 6 Conclusion
In this work, we introduced a compression method for ASR encoders that leverages the inherent low-rank structure of activations in linear layers. By applying the PCA algorithm, this method approximates linear layers with a chain of low-rank matrix multiplications and optimizes self-attention to operate in reduced dimensionality. Our comprehensive evaluation demonstrates that our method achieves a Pareto frontier of accuracy and efficiency, paving the way for more efficient ASR deployments for both on-device and data center environments.